suppose the experiment is formed 100 times and the frequency for each outcome that turned up are as follows:
outcome: HHH,HHT,HTH,HTT,THH,THT,TTH,TTT
fequency: HHH(11), HHT(5), HTH(13), HTT(9), THH(17), THT(15), TTH(18), TTT(12)
what is the probability of getting two heads?

Answer by stanbon(75887)

suppose the experiment is formed 100 times and the frequency for each outcome that turned up are as follows:
outcome: HHH,HHT,HTH,HTT,THH,THT,TTH,TTT
fequency: HHH(11), HHT(5), HTH(13), HTT(9), THH(17), THT(15), TTH(18), TTT(12)
what is the probability of getting two heads?
-------
empirical probability = (5+13+17)/100 = 35/100
-------
theoretical probability = (3/8)
===============
Cheers,
Stan H.